1) sin65 + sin43 – sin29 - sin7=?
Angles are in degrees.
​2) sin(pi cot theta)= cos(pi tan theta)
then cot2theta=?
3)cosxcosysin(x-y) + cosycoszsin(y-z) +    coszcosxsin(z-x) +                 sin(x-y)sin(y-z)sin(z-x)=?

Grade:11

1 Answers

Ravi
9 years ago
2
\sin({\pi} cot \theta)= cos({\pi} tan \theta)

\sin({\pi} cot \theta)= sin({\pi}/2+{\pi} tan \theta)

({\pi} cot \theta)=({\pi}/2+{\pi} tan \theta)

Solve for tan(theta) and find the value of tan2(theta). Reciprocate to get to the value of cot2(theta)
Requested to post only one question per post.